If the cost of 7 m of cloth is ₹ 1470, find the cost of 5 m of cloth. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
Cost of 7 m cloth =₹ 1470
Cost of 1 m cloth = 1470 / 7
= ₹ 210
So, cost of 5 cloth = 210 × 5 = 1050
∴Cost of 5 m cloth is₹ 1050

Thanks